OpenStudy (anonymous):

A ball is thrown vertically upward from a height of 6 feet with an initial velocity of 60 feet for second. How high will the ball go? a(t)=-32 feet per second is the acceleration of gravity.Problem involves integrals.

OpenStudy (anonymous):

\[h(t)=-16t^2+v_0t+h_0\]in your case you have \(v_0=60, h_0=6\) i guess giving you the equation \[h(t)=-16t^2+60t+6\] max is at the vertex, where \(t=-\frac{b}{2a}=-\frac{60}{2\times -16}=\frac{15}{8}\)

OpenStudy (anonymous):

replace \(t\) by \(\frac{15}{8}\) to get the maximum height

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

where did you get the -16 from?

OpenStudy (anonymous):

aah good question first of because i know it. second because the acceleration is constatn at -32 this is the second derivative of the position function so first derivative is \[-32t+v_0\] and position funciton itself if \[-16t^2+v_0t+h_0\] by taking anti derivatives twice

OpenStudy (anonymous):

i should say acceleration due to gravity is -32 feet per second squared

OpenStudy (anonymous):

ohhhok..got it! thanks!!

OpenStudy (anonymous):

yw